The molarity of a NaOH solution was done 3 times. What is the % deviation? You must show 4 distinct steps.

Molarity of NaOH:

0.1034

0.1098

0.1067
Answer #1

Inorder to calculater the % deviation the four steps are required

i) Mean Of data points

ii) Calculation of average deviation from each data point

iii) Average deviation

iv) % deviation = (Average deviation / Mean Of data points ) * 100

The following table gives the result followed by calculation

moles of NaOH mean deviation avg deviation % deviation
0.1034 0.1066 -0.0032 0.002166667 2.032520325
0.1098 0.0032
0.1067 0.0001

Mean = (0.1034 + 0.1098 + 0.1067 ) / 3 = 0.1066

deviation = I 0.1034 - 0.1066 I = 0.0032

= I 0.1098 - 0.1066 I = 0.0032   

= I 0.1067 - 0.1066 I = 0.0001

Average Deviation = (0.0032 + 0.0032 + 0.0001 ) / 3 = 0.002166667

% deviation = ( 0.002166667 / 0.1066 ) * 100 =2.03 %

The  % deviation was found to be 2.03 %